The original arc reactor designed by Tony Stark to power his Iron Man suit was fueled by Pd-103, which is an isotope of palladium metal that decays by electron capture into stable Rh-103. The half-life of Pd-103 is 17 days. A new fuel cell for the arc reactor contains 1.6 grams of Pd-103. Once the amount of Pd-103 in the reactor has fallen to 0.82 grams, it cannot produce sufficient power to run the Iron Man suit. How often does Tony have to replace the palladium fuel cell? Even though Rh-103 is not radioactive, palladium toxicity in Tony’s blood would eventually force him to discover a new element to power the reactor.

Group of answer choices 12.0 days 36.0 days 4.00 days 16.4 days 21.0 days
Answer #1

xya = 17 days lin a ena 17 -> X = 0.693 π 0.040 764.7 17 N = 1-6 gram 0.82 gram N- Now, N= Nee-it Oooyo7647t - 0.82 - lobe. .Hope it helps

Kindly like the answer and comment your doubts in the comment section below.